OpenStudy (anonymous):

Line, parabola, cirlce, ellipse, or hyperbola? 1] 2y=x^2 2] 108y^2-81x=0 3] y=2x+5 4] 10x^2+10y^2=40

OpenStudy (anonymous):

1. parabola 2. parabola 3. line 4. ellipse

OpenStudy (anonymous):

what about: 121x^2+81Y^2=9801 & 121x^2-81y^2=9801

OpenStudy (anonymous):

ellipse hyperbola

OpenStudy (anonymous):

Thank you . (:

OpenStudy (anonymous):

Just ask!

OpenStudy (anonymous):

If you had to put 25x^2-4y^2=100 in standard form, would it be 25x^2-4y^2-100=0?

OpenStudy (anonymous):

No, standard form is the one has radius to the right side!

OpenStudy (anonymous):

So it's already in standard form!

OpenStudy (anonymous):

ok, what would be its center and vertices?

OpenStudy (anonymous):

center (0,0)

OpenStudy (anonymous):

circle just have center and radius!

OpenStudy (anonymous):

ok thanks, what about this: write an equation of the hyperbola with the given foci & vertices: foci: (8,0),(-8,0) Vertices: (7,0),(-7,0)

OpenStudy (anonymous):

Not sure if I 'm correctly, so I'll get back to you with it!